# Work extraction with feedback control using limited resources

### v1

@@ II.2 Introducing control limitations
In contrast to the ideal work extraction scenario described above, in many real-world scenarios the set of possible quenching Hamiltonians is finite and much smaller than the set of possible outcomes of the measurement. In many others that set is predefined and fixed ahead of time, and the only thing one can do is choose one of those possibilities. We are interested in quantifying how much the ideal scenario presented above is affected when introducing such limitations affecting the control abilities of the agent performing work extraction.
We focus in a broad range of scenarios characterized by the impossibility of the controller to perform the ideal quench introduced in Eq. (5), that leads to a post-quench Hamiltonian $H_{\mathbf{m}}(\mathbf{x})=-k_{B}T\ln\rho_{X|\mathbf{m}}(\mathbf{x})$. Instead, we consider situations in which the feedback controller action is limited to choose one out of some finite set of quench Hamiltonians. We write that set as :=$\mathcal{H}:=\{H_{k}(\mathbf{x})\}_{k=1}^{N}$ for the $N$ possible options. In this situation, after reading the measurement results $\mathbf{m}$, the feedback controller selects and applies one of the $N$ quenches in the repertoire $\mathcal{H}$ so as to maximize the amount of extractable work, according to the obtained measurement result. We denote the index leading to the (optimal) choice of quenching Hamiltonian for a given measurement result $\mathbf{m}$ by $n(\mathbf{m})$, to be determined later. We will first assume the form of the elements $H_{k}(\mathbf{x})$ in $\mathcal{H}$ to be predefined and fixed, while in following sections we will also determine their optimal form.
In any case, as a consequence of the limitations in the available actions of the controller for the implementable (post-quench) Hamiltonians, the state of the system after the quench will no longer be in equilibrium with the post-quench Hamiltonian, i.e., $\rho_{X|\mathbf{m}}(\mathbf{x})\neq k_{B}T\ln H_{n}(\mathbf{x})$ for any $n$, and hence an irreversible relaxation from the initial state $\rho_{X|\mathbf{m}}(\mathbf{x})$ to the actual equilibrium state after the quench $e^{-\beta H_{n}(\mathbf{x})}/Z_{n}$ (with $Z_{n}=\sum_{x}e^{-\beta H_{n}(\mathbf{x})}$) will unavoidably take place before the quasistatic process completing the work extraction protocol is performed (or completed). In this case the work extractable from the initial state is reduced and reads:
$W_{n}(\mathbf{m})$ $=-\langle H_{n}\rangle_{\rho_{X|\mathbf{m}}}+\langle H\rangle_{\rho_{X|\mathbf{m}}}+F_{H_{n}}^{\rm eq}-F_{H}^{\rm eq}$ (8)
@@
In the last equation, the first two terms account for the free energy extractable as work in the ideal case (c.f. Eq. (II.1)), where exact quenching to $H_{\mathbf{m}}(\mathbf{x})=-k_{B}T\ln\rho_{X|\mathbf{m}}(\mathbf{x})$ is possible (and hence $Z_{\mathbf{m}}=1$). The second term corresponds to the entropy production (or dissipation) generated in the irreversible post-quench relaxation $\rho_{X|\mathbf{m}}(\mathbf{x})\rightarrow e^{-\beta{H}_{n}(\mathbf{x})}/Z_{n}$, which we can write as
@@
We can hence rewrite Eq. (8) as $W_{n}(\mathbf{m})=W_{\mathrm{ext}}(\mathbf{m})-k_{B}TS_{\mathrm{tot}}(\mathbf{m})$. It is worth noticing that the first term, $W_{\mathrm{ext}}(\mathbf{m})$, does not depend on the actual quenching Hamiltonian chosen in the imperfect scenario $H_{n}(\mathbf{x})$. It is only the second, entropy production term $S_{\mathrm{tot}}(\mathbf{m})$ that depends on that Hamiltonian. Note in particular that if the two distributions in that second term are very different (as quantified KL-divergence), entropy production may be large enough that the net work extracted in the process is negative. More explicitly, that will be the case whenever
@@
Repeating this entire process many times and taking the average of the work extracted in Eq. (8) over the different measurement results, we obtain the average work extracted:
$\langle{W}_{n}\rangle=$ $~{}k_{B}TI(\mathbf{m};\mathbf{x})+\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq}$
$-k_{B}T\langle S_{\mathrm{tot}}\rangle,$ (11)
where the first line gives the work extracted in the ideal case (c.f. Eq. (II.1)) and the second line gives the average entropy production due to imperfect quenching, which reduces the amount of work extracted. That average entropy production over measurement results is given by a a related calculation:
@@
These two calculations will be the basis of our main results. Again, erasure of the information employed for the feedback control in preparation for a next iteration of the entire process will have an unavoidable cost of $k_{B}TI(\mathbf{m};\mathbf{x})$ imposed by the generalized Landauer’s bound.
@@ III The value of information under control limitations
To make things more precise, we start by considering a partition of the entire space $\mathbb{M}^{d_{M}}$ of measurement results $\mathbf{m}$ into $N$ disjoint sets $l(n)$ with $n=1,...,N$, such that $\cup_{n}\ell(n)=\mathbb{M}^{d_{M}}$ and $\ell(n)\cap\ell(n^{\prime})=\varnothing$ for $n\neq n^{\prime}$. In that way, every measurement result $\mathbf{m}$ is unequivocally associated to a set $l(n)$ with label $n$. The probability that the measurement outcome $\mathbf{m}$ is inside the set $l(n)$ can then be written as:
@@
Notice that we by this point we haven’t impose any property on the size of the different sets $l(n)$.
We then ascribe one of the possible quenching Hamiltonians $H_{n}(\mathbf{x})\in\mathcal{H}$ to each set $l(n)$. Since we are considering a finite partition, the computation of $l(n)$ that associates the Hamiltonians in the set $\mathcal{H}$ with the labels $n$ (linked to the observations $\mathbf{m}$) is just a lookup table. This table returns the Hamiltonian $H_{n}(\mathbf{x})$ from the repertoire $\mathcal{H}$ verifying that $e^{-\beta H_{n}(\mathbf{x})}/Z_{n}$ is most similar to $\rho_{X|\mathbf{m}}(\mathbf{x})$ as quantified by Kullback-Leibler (KL) divergence. Formally, the optimal lookup table is the function
@@
In this way the labels $n$ associated to the Hamiltonians $H_{n}(\mathbf{x})$ in $\mathcal{H}$ are determined by the minimization of the entropy production for given outcome $\mathbf{m}$ in Eq. (9) over all the possible entries in $\mathcal{H}$. In particular, the optimal value of the net extracted work given by Eq. (II.2) is
$\langle{W}_{n^{*}}\rangle=$ $~{}k_{B}TI(\mathbf{m};\mathbf{x})+\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq}$
$-k_{B}T\langle S^{*}_{\mathrm{tot}}\rangle,$ (15)
where Eq. 12 gives
$\langle S^{*}_{\mathrm{tot}}\rangle=\sum_{\mathbf{m}}p(\mathbf{m})D\left(\rho_{X|\mathbf{m}}\left|\left|\frac{e^{-\beta H_{n^{*}(\mathbf{m})}}}{Z_{n^{*}(\mathbf{m})}}\right.\right.\right),$ (16)
@@ III.1 Conditions for effective work extraction under control limitations
Having specified the way to choose the different quenching Hamiltonians $H_{n}(\mathbf{x})$ in the set $\mathcal{H}$ for a given measurement outcome $\mathbf{m}$, we may implement feedback control to extract the average amount of work given in Eq. (II.2). That implementation would make sense whenever a net work can be extracted from $\rho_{0}$ once the cost of information erasure is discounted. The requirement for successful feedback control is hence $k_{B}T\langle S_{\mathrm{tot}}\rangle<\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq}$, which can be conveniently written as:
$\langle F_{H_{n}}^{\rm eq}\rangle-F_{H}^{\rm eq}>\langle W_{\rm cost}\rangle+k_{B}TI(\mathbf{m}:\mathbf{x}),$ (17)
@@
$\langle W_{\rm cost}\rangle:=\sum_{x}\rho_{0}(\mathbf{x})\left[\sum_{n}P(n|\mathbf{x})H_{n}(\mathbf{x})-H(\mathbf{x})\right],$ (18)
with $P(n|\mathbf{x})=\sum_{\mathbf{m}\in\ell(n)}P(\mathbf{m}|\mathbf{x})$ being the effect of the measurement noise on the choice of the quenching Hamiltonian $H_{n}$. Equation (17) implies that a net work extraction is possible in the feedback scenario if on average the (equilibrium) free energy difference between the quenching Hamiltonians and the original Hamiltonian $H$, overcomes the combined work cost of implementing the quench and posterior erasure of the information employed in the feedback control.
@@ III.2 Work extraction from deterministic and random quenching operations
Now consider a different scenario in which we try to extract work by using a either a pre-specified or a random quenching Hamiltonian $H_{q}\in\mathcal{H}$ which is actually independent of the measurement result $\mathbf{m}$. If the quenching Hamilronian is fixed to $H_{q}$, the average entropy production in Eq. (12) simplifies to $k_{B}T\langle S_{\mathrm{tot}}\rangle=k_{B}T\sum_{\mathbf{m}}p(\mathbf{m})D(\rho_{X|\mathbf{m}}||e^{-\beta H_{q}}/Z_{q})=k_{B}TI(\mathbf{m};\mathbf{x})+k_{B}TD(\rho_{0}||e^{-\beta H_{q}})$, and hence the work that can be extracted in this case becomes:
$\langle{W}_{q}\rangle$ $=\mathcal{F}_{H}(\rho_{0})-{F}_{H}^{\rm eq}-k_{B}TD\left(\rho_{0}\left|\left|\frac{e^{-\beta H_{q}}}{Z_{q}}\right.\right.\right)$
$=F_{H_{q}}^{\rm eq}-F_{H}^{\rm eq}-\sum_{x}\rho_{0}(x)[H_{q}(\mathbf{x})-H(\mathbf{x})],$ (19)
@@
$W_{\rm cost}^{q}=\sum_{x}\rho_{0}(x)[H_{q}(\mathbf{x})-H(\mathbf{x})],$ (20)
@@
We note that above, if we assume $H_{q}=-\ln\rho_{0}$ (and hence $Z_{q}=1$), the third therm in the first line of Eq. (III.2) vanishes and we can extract the full free energy contained in $\rho_{0}$, namely, $W_{q}=\mathcal{F}_{H}(\rho_{0})-{F}_{H}^{\rm eq}$ in accordance to optimal work extraction without feedback control Hasegawa et al. (2010); Esposito and den Broeck (2011). However, the Hamiltonians in the set $\mathcal{H}$ would be typically optimized to extract work from a specific measurement outcome ${m}$, and hence we assume in the following that $H_{0}\notin\mathcal{H}$.
@@
Comparing the work extractable from imperfect feedback control (once the costs of information erasure has been subtracted) with $\langle\bar{W}_{q}\rangle$ we obtain the work gain due to using information in the quenching as:
$\langle\Delta W\rangle:=\langle{W}_{n}\rangle-k_{B}TI(\mathbf{m};\mathbf{x})-\langle{W}_{q}\rangle$ (21)
$=\sum_{n\neq q}~{}p(n)~{}F_{H_{n}}^{\rm eq}+\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})\left[H_{q}(\mathbf{x})-\sum_{n}P(n|\mathbf{x})H_{n}(\mathbf{x})\right]$
@@
where we used that $H_{q}\in\mathcal{H}$. The first sum of equilibrium free energies in the above expression is always positive. However, $\langle\Delta W\rangle$ can in principle take any sign depending on how $H_{q}$ and the set $\mathcal{H}$ is chosen, as well as on the level of noise in the measurement as captured by $P(n|\mathbf{x})$. In particular, a net benefit in work extraction for the feedback control scenario requires that the energy after the quench $q$ is not very low compared with the average after-quenching energy. This will always be the case in the situation in which the work cost of quenching is negligible (that is, $\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})H\simeq\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})H_{q}$ for any $q$) and hence the second line in the above equation vanishes leading to $\Delta W\geq 0$.
@@
The expression in Eq. (21) hence implies that using feedback control may allow the feedback controller not only to gain an extra amount of work $k_{B}TI(\mathbf{m};\mathbf{x})$ related to the information acquisition from the measurements (which needs to be payed back at the end of the protocol), but it may also increase our ability to extract work from the original distribution $\rho_{0}$ (whenever the optimal quench $H_{0}:=-k_{B}T\ln\rho_{0}$ is not available), enhancing the amount of work we can obtain with respect to the case of just using a predefined quench in $\mathcal{H}$.
@@
If we now assume that the Hamiltonian $H_{q}$ is sampled randomly from a distribution $r(q)$ (independent from the measurement result), the average work extractable would be $\langle\bar{W}\rangle_{r}=\sum_{q}r(q)\langle W_{q}\rangle$ with $\langle W_{q}\rangle$ in Eq. (III.2). In that case the difference between the work extracted when choosing the correct quench and a random one from $r(q)$ is:
$\langle\Delta\bar{W}\rangle_{r}$ $=\langle{W}_{n}\rangle-k_{B}TI(\mathbf{m};\mathbf{x})-\sum_{q}r(q)\langle{W}_{q}\rangle$
$=\sum_{n}~{}[p(n)-r(n)]~{}F_{H_{n}}^{\rm eq}$ (22)
$+\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})\sum_{n}[r(n)-P(n|\mathbf{x})]H_{n}(\mathbf{x}).$
@@
We note that if we choose the random distribution above to mimic the marginal distribution to obtain the measurement outcomes, $r(q)=p(q)$, the first term in the above equation vanishes. Moreover, since $p(n)=\sum_{x}P(n|\mathbf{x})\rho_{0}(\mathbf{x})$ the second term is positive and we have a net gain $\langle\Delta\bar{W}\rangle_{p}\geq 0$.

### v2

@@ II.2 Introducing control limitations
In contrast to the ideal work extraction scenario described above, in many real-world scenarios the set of possible quenching Hamiltonians is finite and much smaller than the set of possible outcomes of the measurement. In addition, often in the real world that set is predefined and fixed ahead of time, and the only thing one can do is choose one of those fixed possibilities. We are interested in quantifying how much the ideal scenario presented above is affected when introducing such limitations affecting the control abilities of the agent performing work extraction.
More specifically, we consider a broad range of real-world scenarios characterized by the impossibility of the controller to perform the ideal quench introduced in Eq. (5), that leads to a post-quench Hamiltonian $H_{\mathbf{m}}(\mathbf{x})=-k_{B}T\ln\rho_{X|\mathbf{m}}(\mathbf{x})$. Instead, we consider the situation in which the feedback controller action is limited to choose one out of some finite set of quench Hamiltonians. We write that set as :=$\mathcal{H}:=\{H_{k}(\mathbf{x})\}_{k=1}^{N}$ with $N<|\mathbb{M}^{d_{M}}|$ possible options. In this situation, after reading the measurement results $\mathbf{m}$, the feedback controller selects and applies one of the quenches in the repertoire $\mathcal{H}$ so as to maximize the amount of extractable work, according to the obtained measurement result. We denote the choice by the control protocol of which Hamiltonian to quench to in response to a given measurement result $\mathbf{m}$ by $n(\mathbf{m})$, or $n$ for short. (We derive the optimal form of that function below.) Note that because $\mathbf{m}$ is random, in general so will $n$ be. We will first assume the form of the elements $H_{k}(\mathbf{x})$ in $\mathcal{H}$ to be predefined and fixed, while in following sections we will also determine their optimal form.
In any case, as a consequence of the limitations in the available actions of the controller for the implementable (post-quench) Hamiltonians, the state of the system after the quench will no longer be in equilibrium with the post-quench Hamiltonian. In other words, $\rho_{X|\mathbf{m}}(\mathbf{x})\neq-k_{B}T\ln H_{n(\mathbb{m})}(\mathbf{x})$.
Hence an irreversible relaxation from the (initial) state $\rho_{X|\mathbf{m}}(\mathbf{x})$ to the actual equilibrium state after the quench $e^{-\beta H_{n}(\mathbf{x})}/Z_{n}$ (with $Z_{n}=\sum_{x}e^{-\beta H_{n}(\mathbf{x})}$) will unavoidably take place before the quasi-static process completing the work extraction protocol is completed. In this case the work extractable from the initial state using feedback control is reduced:
$W(\mathbf{m})$ $=-\langle H_{n}\rangle_{\rho_{X|\mathbf{m}}}+\langle H\rangle_{\rho_{X|\mathbf{m}}}+F_{H_{n}}^{\rm eq}-F_{H}^{\rm eq}$ (8)
@@
In the last equation, the first two terms account for the free energy extractable as work in the ideal case [c.f. Eq. (II.1)], where exact quenching to $H_{\mathbf{m}}(\mathbf{x})=-k_{B}T\ln\rho_{X|\mathbf{m}}(\mathbf{x})$ is possible (and hence $Z_{\mathbf{m}}=1$). The second term corresponds to the entropy production (or dissipation) generated in the irreversible post-quench relaxation $\rho_{X|\mathbf{m}}(\mathbf{x})\rightarrow e^{-\beta{H}_{n}(\mathbf{x})}/Z_{n}$, which we can write as
@@
We can hence rewrite Eq. (8) as $W(\mathbf{m})=W_{\mathrm{ideal}}(\mathbf{m})-k_{B}TS_{\mathrm{tot}}(\mathbf{m})$. It is worth noticing that the first term, $W_{\mathrm{ideal}}(\mathbf{m})$, does not depend on the actual quenching Hamiltonian chosen in the imperfect scenario $H_{n}(\mathbf{x})$. It is only the second, entropy production term $S_{\mathrm{tot}}(\mathbf{m})$ that depends on that Hamiltonian. Note in particular that if the two distributions in that second term are very different (as quantified by their KL-divergence), entropy production may be large enough that the net work extracted in the process is negative. More explicitly, that will be the case whenever
@@
Repeating this entire process many times and taking the average of the work extracted in Eq. (8) over the different measurement results, we obtain the average work extracted under control limitations:
$\langle{W}\rangle=$ $~{}k_{B}TI_{X;M}+\Delta\mathcal{F}_{H}-k_{B}T\langle S_{\mathrm{tot}}\rangle,$ (11)
where we used the short-hand notation $\Delta\mathcal{F}_{H}:=\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq}$. In the above equation the first two terms give the work extracted in the ideal case [c.f. Eq. (II.1)] and the third one gives the average entropy production due to imperfect quenching, which reduces the amount of work extracted. That average entropy production over measurement results is given by a a related calculation:
@@
These two calculations will be the basis of our main results. Again, erasure of the information employed for the feedback control in preparation for a next iteration of the entire process will have an unavoidable cost of $k_{B}TI_{X;M}$ imposed by the generalized Landauer’s bound Parrondo et al. (2015).
@@ III The value of information under control limitations
To make things more precise, we start by considering a partition of the entire space $\mathbb{M}^{d_{M}}$ of measurement results $\mathbf{m}$ into $N$ disjoint sets ${L}:=\{l(n)\}_{n}$ with $n=1,...,N$, such that $\cup_{n}\ell(n)=\mathbb{M}^{d_{M}}$ and $\ell(n)\cap\ell(n^{\prime})=\varnothing$ for $n\neq n^{\prime}$. In that way, every measurement result $\mathbf{m}$ can be unequivocally associated to a set $l(n)$ with label $n$. The probability that the measurement outcome $\mathbf{m}$ is inside the set $l(n)$ can then be written as:
@@
Notice that by this point we haven’t impose any property on the size of the different sets $l(n)$.
We then ascribe one of the possible quenching Hamiltonians $H_{n}(\mathbf{x})\in\mathcal{H}$ to each set $l(n)\in{L}$. Since we are considering a finite partition, the computation of $l(n)$ that associates the Hamiltonians in the set $\mathcal{H}$ with the labels $n$ (linked to the observations $\mathbf{m}$) is just a lookup table. This table returns the Hamiltonian $H_{n}(\mathbf{x})$ from the repertoire $\mathcal{H}$ verifying that $e^{-\beta H_{n}(\mathbf{x})}/Z_{n}$ is most similar to $\rho_{X|\mathbf{m}}(\mathbf{x})$ as quantified by Kullback-Leibler (KL) divergence. Formally, the optimal lookup table is the function
@@
In this way the labels $n$ associated to the Hamiltonians $H_{n}(\mathbf{x})$ in $\mathcal{H}$ are determined by the minimization of the entropy production for given outcome $\mathbf{m}$ in Eq. (9) over all the possible entries in $\mathcal{H}$. In particular, the optimal value of the net extracted work is given by application of Eq. (14) in Eqs. (11) and (12).
@@ III.1 Conditions for effective work extraction under control limitations
Having specified the way to choose the different quenching Hamiltonians $H_{n}(\mathbf{x})$ in the set $\mathcal{H}$ for a given measurement outcome $\mathbf{m}$, we may implement feedback control to extract the average amount of work given in Eq. (11). That implementation would make sense whenever a net work can be extracted from $\rho_{0}$ once the cost of information erasure is discounted. The requirement for successful feedback control is hence $k_{B}T\langle S_{\mathrm{tot}}\rangle<\Delta\mathcal{F}_{H}$, which can be conveniently written as:
$\langle F_{H_{n}}^{\rm eq}\rangle-F_{H}^{\rm eq}>\langle W_{\rm cost}\rangle+k_{B}TI_{X;M},$ (15)
@@
$\langle W_{\rm cost}\rangle:=\sum_{x}\rho_{0}(\mathbf{x})\left[\sum_{n}p_{{L}|\mathbb{x}}(n)H_{n}(\mathbf{x})-H(\mathbf{x})\right],$ (16)
with $p_{{L}|\mathbb{x}}(n):=p(n|\mathbf{x})=\sum_{\mathbf{m}\in\ell(n)}p_{M|\mathbb{x}}(\mathbf{m}|\mathbf{x})$ being the effect of the measurement noise on the choice of the partition $l(n)$ corresponding to the quenching Hamiltonian $H_{n}$.Equation (15) implies that a net work extraction is possible using feedback control if on average the (equilibrium) free energy difference between the quenching Hamiltonians and the original Hamiltonian $H$, overcomes the combined work cost of implementing the quench and posterior erasure of the information employed in the feedback control.
@@ III.2 Work extraction from deterministic and random quenching operations
Now consider a different scenario in which we try to extract work by using a either a pre-specified or a random quenching Hamiltonian $H_{q}\in\mathcal{H}$ which is actually independent of the measurement result $\mathbf{m}$. If the quenching Hamilronian is fixed to $H_{q}$, the average entropy production in Eq. (12) simplifies to $\langle S_{\mathrm{tot}}\rangle=\sum_{\mathbf{m}}p(\mathbf{m})D(\rho_{X|\mathbf{m}}||e^{-\beta H_{q}}/Z_{q})=I_{X;M}+D(\rho_{0}||e^{-\beta H_{q}})$, and hence the work that can be extracted in this case becomes:
${W}_{q}$ $=\mathcal{F}_{H}(\rho_{0})-{F}_{H}^{\rm eq}-k_{B}TD\left(\rho_{0}\left|\left|\frac{e^{-\beta H_{q}}}{Z_{q}}\right.\right.\right)$
$=F_{H_{q}}^{\rm eq}-F_{H}^{\rm eq}-W_{\rm cost}^{q},$ (17)
@@
$W_{\rm cost}^{q}=\sum_{x}\rho_{0}(x)[H_{q}(\mathbf{x})-H(\mathbf{x})].$ (18)
@@
We note that above, if we assume $H_{q}=H_{0}:=-k_{B}T\ln\rho_{0}$ (and hence $Z_{q}=1$), the third therm in the first line of Eq. (III.2) vanishes and we can extract the full free energy contained in $\rho_{0}$, namely, $W_{q}=\mathcal{F}_{H}(\rho_{0})-{F}_{H}^{\rm eq}$ in accordance to optimal work extraction without feedback control Takara et al. (2010); Hasegawa et al. (2010); Esposito and den Broeck (2011). However, the Hamiltonians in the set $\mathcal{H}$ would be typically optimized to extract work from a specific measurement outcome ${m}$, and hence we assume in the following that $H_{0}\notin\mathcal{H}$.
@@
Comparing the work extractable from imperfect feedback control (once the costs of information erasure has been subtracted) with ${W}_{q}$ we obtain the work gain due to using information in the quenching as:
$\langle\Delta W\rangle$ $:=\langle{W}\rangle-k_{B}TI_{X;M}-{W}_{q}$
$=\sum_{n\neq q}~{}p(n)~{}F_{H_{n}}^{\rm eq}$ (19)
$~{}+\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})\left[H_{q}(\mathbf{x})-\sum_{n}p_{{L}|\mathbf{x}}(n)H_{n}(\mathbf{x})\right]$
@@
where we used that $H_{q}\in\mathcal{H}$. The first sum of equilibrium free energies in the above expression is always positive. However, $\langle\Delta W\rangle$ can in principle take any sign depending on how $H_{q}$ and the set $\mathcal{H}$ is chosen, as well as on the level of noise in the measurement as captured by $p_{{L}|\mathbb{x}}(n)$. In particular, a net benefit in work extraction for the feedback control scenario requires that the energy after the quench $q$ is not very low compared with the average after-quenching energy. This will always be the case in the situation in which the work cost of quenching is negligible (that is, $\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})H\simeq\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})H_{q}$ for any $q$) and hence the second line in the above equation vanishes leading to $\langle\Delta W\rangle\geq 0$.
@@
The expression in Eq. (III.2) hence implies that using feedback control may allow the feedback controller not only to gain an extra amount of work $k_{B}TI_{X;M}$ related to the information acquisition from the measurements (which needs to be paid back at the end of the protocol), but it may also increase our ability to extract work from the original distribution $\rho_{0}$ (whenever the optimal quench $H_{0}:=-k_{B}T\ln\rho_{0}$ is not available), enhancing the amount of work we can obtain with respect to the case of just using a predefined quench in $\mathcal{H}$.
@@
If we now assume that the Hamiltonian $H_{q}$ is sampled randomly from a distribution $r(q)$ independent from the measurement result, the average work extractable would be $\langle{W}\rangle_{r}=\sum_{q}r(q)W_{q}$ with $W_{q}$ in Eq. (III.2). In that case the difference between the work extracted when choosing the correct quench and a random one from $r(q)$ is:
$\langle\Delta W\rangle_{r}$ $=\langle{W}\rangle-k_{B}TI_{X;M}-\sum_{q}r(q){W}_{q}$
$=\sum_{n}~{}\left[p(n)-r(n)\right]~{}F_{H_{n}}^{\rm eq}$ (20)
$+\sum_{\mathbf{x}}\rho_{0}(\mathbf{x})\sum_{n}\left[r(n)-p_{{L}|\mathbf{x}}(n)\right]H_{n}(\mathbf{x}).$
@@
We note that if we choose the random distribution above to mimic the marginal distribution to obtain the action from the measurement outcomes, $r(q)=p(q)$, the first term in the above equation vanishes. Moreover, since $p(n)=\sum_{x}p_{{L}|\mathbf{x}}(n)\rho_{0}(\mathbf{x})$ the second term is positive and we have a net gain $\langle\Delta W\rangle_{p}\geq 0$, remarking the convenience of using feedback control.